# Random matrix universality in dynamical correlation functions at late times

### v2

@@ 3.2.1 GOE spin glass Hamiltonian
Figure 5: Left: $(a)$ Shifted correlation function $\overline{\mathcal{C}}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 XY spin glass Hamiltonians (46) at inverse temperature $\beta=0$. Right: $(c)$ Shifted correlation function $\mathcal{C}(t)$ and $(d)$ spectral form factor SFF$(t)$ obtained from a single XY spin glass Hamiltonian (46). Plots $(c)$ and $(d)$ are for $L=14$ spins at inverse temperature $\beta=0$. In all plots, the thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@
As in [38], $F^{(2)}(\omega)$ can be computed for the XY spin glass by simply Fourier transforming the shifted correlation function $\mathcal{C}(t)$ shown in Fig. 5. The result shown in Fig. 4(right) checks assumption 2. It is also worth noticing that $F^{(2)}(\omega)$ is well fitted by a sum of two shifted Lorentzians at small frequencies, which means that $\mathcal{C}_{\text{d}}$ displays an exponential decay with oscillations within (as seen in Fig. 5).
The numerical results for the ${\rm SFF}(t)$ and $\mathcal{C}(t)$ are shown for a single Hamiltonian in Fig. 5 (right), and averaging over 1000 Hamiltonians in Fig. 5 (left). The shifted correlation function $\overline{\mathcal{C}}(t)$ exhibits a slope-dip-ramp-plateau behaviour, just as the SFF, very visible in the left plot of Fig. 5. However, for the system sizes at our disposal, the ramp is not as linear as in the SFF, pointing to the effect of corrections in the kernel induced by $f(\omega)$. Moreover, we note that differently from the SFF, the function $\mathcal{C}$ is not positively defined, so in the slope, it shows negative oscillations. The fact that, instead, it remains positive at large times is a consequence of the predictions that we make about its random matrix component. Moreover, as in RMT, we clearly see that, contrarily to the SFF, the time average of $\mathcal{C}$ does not reproduce the ensemble average in the ramp.
@@
3.2.2 GUE spin-class Hamiltonian
@@
Figure 7: Left: $(a)$ Shifted correlation function $\overline{\mathcal{C}}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 chiral spin glass Hamiltonians (47) at inverse temperature $\beta=0$. Right: $(c)$ Shifted correlation function $\mathcal{C}(t)$ and $(d)$ and spectral form factor SFF$(t)$ obtained from a single chiral spin glass Hamiltonian (47). Plots $(c)$ and $(d)$ are for $L=14$ spins at inverse temperature $\beta=0$. In all plots, the thin curves are the exact result, while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@
where $\varepsilon_{lmn}$ is the standard Levi-Civita symbol and $l,m,n$ take the values of all possible permutations of $i,j,k$. From Eq. (48), it follows that $[H,m]=0$ with $m=\frac{1}{L}\sum_{i}S_{i}^{z}$ the magnetization. For an even number $L$ of spins, the analysis can, therefore, be restricted to the magnetization sector $m=2/L$ considering the observable $S^{z}_{L/2}$. Numerically, it appears that this model has a few degenerate levels but, because they are so few, we can forget about them for large systems (i.e. $L$ greater than $\sim 10$). As in the previous subsection, the assumptions made in Sec. 3.1 are fulfilled as seen in Fig. 6. Moreover, the on-shell correlations $F^{(2)}(\omega)$ are well fitted by a Lorentzian at small frequency so $\mathcal{C}_{\text{d}}$ has an exponential decay in time.
The SFF$(t)$ and $\mathcal{C}(t)$ were computed through an exact numerical diagonalization of the Hamiltonian (47). The results are shown for a single Hamiltonian in Fig. 7 (right), and averaging over 1000 Hamiltonians in Fig. 7 (left). The slope-dip-ramp-plateau feature is well visible in this spin glass system, in both $\mathcal{C}(t)$ and SFF$(t)$. Similarly to the previous case, the time average fails to reproduce the ensemble average.
@@ 3.2.3 Characteristic time scales
From Figs. 5 and 7, it appears that for many-body systems, both the ${\rm SFF}$ and the correlation function $\mathcal{C}$ exhibit a similar ramp and plateau. However, the slope appearing at early times is not universal and is expected to differ for the ${\rm SFF}$ and $\mathcal{C}$. Since the dip time $t_{\rm dip}$ signals the crossover between the slope and the ramp, it is not the same for the ${\rm SFF}$ and for $\mathcal{C}$ as seen in Fig. 8. On the contrary, the ramp and the plateau being two universal features, the Heisenberg time $t_{\rm Heis}$, corresponding to the onset of the plateau, is unique.
Figure 8: Dip times $t_{\rm dip}$ and Heisenberg times $t_{\rm Heis}$ of the ${\rm SFF}$ and the correlation function $\mathcal{C}$ as a function of the Hilbert space dimension $\mathcal{N}$. Results for the XY spin glass (46) are in the left plot, and those for the chiral spin glass (47) are on the right. The dip times are extracted at the minimum of the ${\rm SFF}$ (or $\mathcal{C}$) before the ramp. The Heisenberg times are defined as the first time the ${\rm SFF}$ (or $\mathcal{C}$) reaches $95\%$ of its plateau value given by the diagonal ensemble.
@@ 4 Summary and Conclusions
  • In both the RMT and Hamiltonian cases, the symmetry classes are crucial for extracting eigenvalue correlations from properly shifted dynamical correlation functions, defined as $\mathcal{C}(t)=C(t)-(1-\beta_{\rm RMT}/2)\lim_{t\to\infty}\overline{C}(t)$, where $C(t)$ is the connected correlator. This is particularly relevant for the orthogonal ensemble ($\beta_{\rm RMT}=1$), where the ramp is not observed without this shift.
  • While the ensemble average of the SFF always coincides with a small time-window average, the same does not hold true for correlation functions during the ramp phase. This non-self-averaging behaviour of $\mathcal{C}(t)$ emerges from the large relative fluctuations of the matrix elements of the observable under scrutiny.

### v4

@@ 3.2.1 GOE spin glass Hamiltonian
Figure 5: Left: $(a)$ Shifted correlation function $\overline{\Gamma}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 XY spin glass Hamiltonians (46) at inverse temperature $\beta=0$. Right: $(c)$ Shifted correlation function $\Gamma(t)$ and $(d)$ spectral form factor SFF$(t)$ obtained from a single XY spin glass Hamiltonian (46). The inset shows some negative oscillations appearing in the ramp. Plots $(c)$ and $(d)$ are for $L=14$ spins at inverse temperature $\beta=0$. In all plots, the thin curves are the exact result while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@
As in [39], $F^{(2)}(\omega)$ can be computed for the XY spin glass by simply Fourier transforming the shifted correlation function $\Gamma(t)$ shown in Fig. 5. The result shown in Fig. 4(right) checks assumption 2. It is also worth noticing that $F^{(2)}(\omega)$ is well fitted by a sum of two shifted Lorentzians at small frequencies, which means that $\Gamma_{\text{d}}$ displays an exponential decay with oscillations within (as seen in Fig. 5).
The numerical results for the ${\rm SFF}(t)$ and $\Gamma(t)$ are shown for a single Hamiltonian in Fig. 5 (right), and averaging over 1000 Hamiltonians in Fig. 5 (left). The shifted correlation function $\overline{\Gamma}(t)$ exhibits a slope-dip-ramp-plateau behaviour, just as the SFF, very visible in the left plot of Fig. 5. However, for the system sizes at our disposal, the ramp is not as linear as in the SFF, pointing to the effect of corrections in the kernel induced by $f(\omega)$. Moreover, we note that differently from the SFF, the function $\Gamma$ is not positively defined, so in the slope, it shows negative oscillations. The fact that, instead, it remains positive at large times is a consequence of the predictions that we make about its random matrix component. Moreover, as in RMT, we clearly see that, contrarily to the SFF, the time average of $\Gamma$ does not reproduce the ensemble average in the ramp. The noise observed in Fig. 5 could be suppressed introducing some dissipation [40].
@@
3.2.2 GUE spin glass Hamiltonian
@@
Figure 7: Left: $(a)$ Shifted correlation function $\overline{\Gamma}(t)$ and $(b)$ spectral form factor $\overline{{\rm SFF}}(t)$ averaged over 1000 chiral spin glass Hamiltonians (47) at inverse temperature $\beta=0$. Right: $(c)$ Shifted correlation function $\Gamma(t)$ and $(d)$ and spectral form factor SFF$(t)$ obtained from a single chiral spin glass Hamiltonian (47). Plots $(c)$ and $(d)$ are for $L=14$ spins at inverse temperature $\beta=0$. In all plots, the thin curves are the exact result, while the thick curves are a smoothed version obtained by convoluting with a Gaussian kernel. The horizontal dashed lines are the expected plateau values coming from the diagonal ensemble. The increasing oblique dashed lines are fits tproportional-$\propto t$.
@@
where $\varepsilon_{lmn}$ is the standard Levi-Civita symbol and $l,m,n$ take the values of all possible permutations of $i,j,k$. From Eq. (48), it follows that $[H,m]=0$ with $m=\frac{1}{L}\sum_{i}S_{i}^{z}$ the magnetization. For an even number $L$ of spins, the analysis can, therefore, be restricted to the magnetization sector $m=2/L$ considering the observable $S^{z}_{L/2}$. Numerically, it appears that this model has a few degenerate levels but, because they are so few, we can forget about them for large systems (i.e. $L$ greater than $\sim 10$). As in the previous subsection, the assumptions made in Sec. 3.1 are fulfilled as seen in Fig. 6. Moreover, the on-shell correlations $F^{(2)}(\omega)$ are well fitted by a Lorentzian at small frequency so $\Gamma_{\text{d}}$ has an exponential decay in time.
The SFF$(t)$ and $\Gamma(t)$ were computed through an exact numerical diagonalization of the Hamiltonian (47). The results are shown for a single Hamiltonian in Fig. 7 (right), and averaging over 1000 Hamiltonians in Fig. 7 (left). The slope-dip-ramp-plateau feature is well visible in this spin glass system, in both $\Gamma(t)$ and SFF$(t)$. Similarly to the previous case, the time average fails to reproduce the ensemble average.
@@ 3.2.3 Characteristic time scales
From Figs. 5 and 7, it appears that for many-body systems, both the ${\rm SFF}$ and the correlation function $\Gamma$ exhibit a similar ramp and plateau. However, the slope appearing at early times is not universal and is expected to differ for the ${\rm SFF}$ and $\Gamma$. Since the dip time $t_{\rm dip}$ signals the crossover between the slope and the ramp, it is not the same for the ${\rm SFF}$ and for $\Gamma$ as seen in Fig. 8. On the contrary, the ramp and the plateau being two universal features, the Heisenberg time $t_{\rm Heis}$, corresponding to the onset of the plateau, is unique.
Figure 8: Dip times $t_{\rm dip}$ and Heisenberg times $t_{\rm Heis}$ of the ${\rm SFF}$ and the correlation function $\Gamma$ as a function of the Hilbert space dimension $\mathcal{N}$. Results for the XY spin glass (46) are in the left plot, and those for the chiral spin glass (47) are on the right. The dip times are extracted at the minimum of the ${\rm SFF}$ (or $\Gamma$) before the ramp. The Heisenberg times are defined as the first time the ${\rm SFF}$ (or $\Gamma$) reaches $95\%$ of its plateau value given by the diagonal ensemble.
@@ 4 Summary and Conclusions
  • In both the RMT and Hamiltonian cases, the symmetry classes are crucial for extracting eigenvalue correlations from properly shifted dynamical correlation functions, defined as $\Gamma(t)=C(t)-(1-\beta_{\rm RMT}/2)\lim_{t\to\infty}\overline{C}(t)$, where $C(t)$ is the connected correlator. This is particularly relevant for the orthogonal ensemble ($\beta_{\rm RMT}=1$), where the ramp is not observed without this shift.
  • While the ensemble average of the SFF always coincides with a small time-window average, the same does not hold true for correlation functions during the ramp phase. This non-self-averaging behaviour of $\Gamma(t)$ emerges from the large relative fluctuations of the matrix elements of the observable under scrutiny.